Questions 2-8 refer to the following: You and your 60 kg workout partner, Kim, are tossing a 4.0 kg medicine ball back and forth to each other. Shortly before she releases the ball, Kim is in the air (her feet are off the ground). At this instant, the only forces she exerts against the ball are a 200 N horizontal force directed forwards and a 250 N force directed upwards. These forces and gravity are the only forces acting on the ball. The reactions to these forces are the only forces, other than gravity, acting on Kim. (6 points for each question)

6. If the ball has a downward velocity of 4 m/s when you begin to catch it, what total vertical impulse must act on the ball in order to stop its downward motion?

7. What average vertical force must you exert on the ball in order to stop its downward motion in 0.80 second?

8.If you continue to exert this average vertical force on the ball for 2.2 more seconds (for a total force duration of 3.0 seconds) until you release the ball, what is the ball's vertical velocity at release?
Answer #1

Q.6 i mpulse = change in momentum = mu - mv = 60 (4) - 60 (0) = 240 kg m/s vertical impulse must be imparted on the ball to stop it

Q. 7 : forc e= impulse x tim e= 240 ( 0.8) = 192 N

Q.8: force x time = mv - mu

192 ( 2.2) = 6 0 ( V)

v( velocity at release) = 7.04 m/s apprx